OpenStudy (anonymous):

Which point lies on the graph of the following piecewise function? y = x^2 if x <= 0 y = x + 1 if x > 0

OpenStudy (anonymous):

need the graph

OpenStudy (anonymous):

(–1, 0) (1, 1) (0, 1) (–2, 4) Those are the options but there is no graph

OpenStudy (anonymous):

even if you can help me know how to know which point lie on the graph that would be a big help

OpenStudy (anonymous):

ahh wait you can draw it

OpenStudy (anonymous):



OpenStudy (anonymous):

that is your piecewise function

OpenStudy (anonymous):

(-1,0) does not lie o your graph also if x<=0 you use x^2=f(x)=y so 0=1 which is false

OpenStudy (anonymous):

so im guessing (0,1) or (-2,4)

OpenStudy (anonymous):

next (1,1) 1=2 not true

OpenStudy (anonymous):

(0,1) since it is equal to 0 you use the first equation 1=0^2 which is 0, so not true

OpenStudy (anonymous):

lastly (-2,4) it's less than 0 use first equations 4=-2^2=4

OpenStudy (anonymous):

(-2,4

OpenStudy (anonymous):

wow this is easier than im making it....i dont know what i was trying to do but you made it much easier i can actually do the others myself know thank you!

OpenStudy (anonymous):

*now